OpenStudy (anonymous):

Find two positive numbers whose sum is 100 and the sum of whose squares is a minimum. Please and thank youuu!

jimthompson5910 (jim_thompson5910):

"two positive numbers whose sum is 100" ---> x+y = 100 x+y = 100 y = 100-x ------------------------------------------ "the sum of whose squares is a minimum" z = x^2 + y^2 z = x^2 + (100-x)^2 z = x^2 + 10000 - 200x + x^2 z = 2x^2 - 200x + 10000 So you'll then want to find the vertex of 2x^2 - 200x + 10000

OpenStudy (anonymous):

thank you!!

jimthompson5910 (jim_thompson5910):

yw